# ADEdgeDrop: Adversarial Edge Dropping for Robust Graph Neural Networks

### v1

@@ Abstract
Although Graph Neural Networks (GNNs) have exhibited the powerful ability to gather graph-structured information from neighborhood nodes via various message-passing mechanisms, the performance of GNNs is limited by poor generalization and fragile robustness caused by noisy and redundant graph data. As a prominent solution, Graph Augmentation Learning (GAL) has recently received increasing attention. Among prior GAL approaches, edge-dropping methods that randomly remove edges from a graph during training are effective techniques to improve the robustness of GNNs. However, randomly dropping edges often results in bypassing critical edges, consequently weakening the effectiveness of message passing. In this paper, we propose a novel adversarial edge-dropping method (ADEdgeDrop) that leverages an adversarial edge predictor guiding the removal of edges, which can be flexibly incorporated into diverse GNN backbones. Employing an adversarial training framework, the edge predictor utilizes the line graph transformed from the original graph to estimate the edges to be dropped, which improves the interpretability of the edge-dropping method. The proposed ADEdgeDrop is optimized alternately by stochastic gradient descent and projected gradient descent. Comprehensive experiments on six graph benchmark datasets demonstrate that the proposed ADEdgeDrop outperforms state-of-the-art baselines across various GNN backbones, demonstrating improved generalization and robustness.
@@
Graph convolutional network, over-smoothing, semi-supervised classification, graph representation learning.
@@ I Introduction
As powerful data representations, graphs can represent data items and their complex relations by nodes and edges, respectively. To depict node attributes and rich relational features of edges, diverse Graph Neural Networks (GNNs) have been proposed to extract distinctive node and edge representations [1, 2]. Representative GNN architectures include Graph Convolutional Network (GCN) [3], Graph SAmple and aggreGatE (GraphSAGE) [4], Graph Attention Network (GAT) [5], Simplifying Graph Convolution (SGC) [6], etc. During network training, GNNs propagate node features as messages, which are then passed through edges and aggregated [7]. However, inefficient message passing caused by low-quality graph data with noise and redundancy often lets GNNs suffer from overfitting, poor generalization, and less robustness [8, 9, 10].
@@
To enhance the generalization and robustness of GNN models, Graph Augmentation Learning (GAL) has emerged as a powerful technique for improving node representations through diverse data augmentation methods. Different from data augmentation in computer vision, which generates rotated or cropped images to enrich datasets, GAL focuses on learning new graph structures or updating node attributes with a diversity of strategies. One well-known GAL method is utilizing node feature perturbations with random corruption or an adversarial training process, thereby leading to more generalized and robust graph representations [11, 12]. Another popular approach in GALs is edge dropping, which removes or inhibits redundant edges in a graph during training to promote the robustness of GNNs. This process is analogous to the dropout technique used for training deep neural networks to mitigate overfitting. As the most commonly used edge-dropping method, the random edge-dropping strategy [13] removes edges with a certain probability, as illustrated in the left part of Figure 1. However, randomly dropping edges often results in the neglect of critical connections during training as indicated in [14, 15, 16]. Meanwhile, a graph may include noisy edges or redundant edges, which can be replaced by the integration of other multi-hop paths. Therefore, it is crucial to remove those insignificant edges while retaining those critical edges, thereby motivating us to design a new edge-dropping technique for GAL.
@@
To cope with the aforementioned issues, we propose a novel edge-dropping method, namely ADversarial Edge Drop (ADEdgeDrop) for arbitrary GNNs. Specifically, we design an adversarial edge predictor, which is responsible for determining whether to remove an edge according to the estimated edge embedding trained with a line graph derived from the original graph, as illustrated in the right part of Figure 1. To our best knowledge, this is the first endeavor to design a novel supervised edge-dropping method, which leverages an adversarial edge predictor to effectively remove insignificant edges while preserving critical connections. Compared with random edge-dropping methods, the proposed method has the following advantages. (1) Better semantic interpretability owing to the edge predictor that evaluates the dropping probability of each edge through the line graph (derived from the original graph), which effectively captures the relationships between the original edges. This process is supervised by the node attribute homogeneity information that evaluates the probability of correlation between two nodes. (2) Improved robustness and generalization owing to the adversarial training, which can alleviate the overfitting effect of the edge predictor and the downstream classifier. Different from existing methods that perform adversarial training on node features, the proposed ADEdgeDrop learns a more robust graph structure by the adversarial training on edge embedding in a bottom-up manner.
In summary, the technical contributions of this paper are summarized as follows.
@@
  • Comprehensive experiments validate that the proposed method obtains superior performance than existing methods with representative backbone GNNs on six different benchmark datasets by improving the robustness and generalization of the learned graph structure.
@@ II-A Graph Neural Networks
As a powerful tool for handling graph-structured data, GNNs have garnered significant attention, which play a crucial role in graph representation learning by effectively gathering neighborhood features and propagating vital information across the graph. As a well-known example of GNNs, GCNs simplify the graph convolution operator on the non-Euclidean spaces by utilizing a truncated approximation of Chebyshev polynomials [3]. Additionally, Velickovic et al. defined a graph attention mechanism to explore the implicit assignment of neighborhood weights for each node [5]. Moreover, Hamilton et al. proposed GraphSAGE, a method that learns node embeddings by sampling and accumulating representations from local neighborhood nodes [4]. Furthermore, recent research attempts have been made to design simplified GNN variants to enhance GNN computations [6, 17, 18]. With the rapid development of GNNs, considerable studies have emphasized the significance of graph augmentation to enhance the GNN training and promote the robustness.
@@ II-B Graph Augmentation Learning
GAL aims to enhance the graph structure through data augmentation strategies, which are beneficial in addressing various robustness issues of GNNs [19]. On one hand, some endeavors have adopted adversarial training for graph representation to achieve robust node embeddings [20, 21, 22]. On the other hand, graph structure learning methods have attempted to learn richer connections from the original sparse graph to augment the training data. For example, Zhao et al. proposed the GAUG method to explore likely missing edges and remove existing edges by an edge predictor [23] though it may encounter huge computational costs, particularly when dealing with a large number of nodes. As a result, edge-dropping or node-dropping strategies on existing connections are economical approaches to mitigate computational costs since recent studies have demonstrated the promising performance of these dropping methods [24], such as DropEdge [13] and DropNode [25]. Additionally, Fang et al. devised a general message-dropping method that simultaneously considers various types of dropping strategies [26]. However, the quality of a graph is closely related to its node relationships, thereby directly influencing the performance of downstream tasks. Randomly removing edges without any criteria may have a negative impact on the graph representation, particularly if important connections are frequently eliminated during training. Hence, the key challenge of edge-dropping methods lies in identifying and removing incongruous connections rather than the essential node relationships.
@@ III-A Notations
In this paper, we define a graph as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ is the set of vertices and $\mathcal{E}$ is the set of edges. The feature matrix and the adjacency matrix of the graph $\mathcal{G}$ are denoted by $\mathbf{X}$ and $\mathbf{A}$, respectively. Similarly, the line graph transformed from the original graph $\mathcal{G}$ is represented as $\mathcal{G}_{lg}=\{\mathcal{V}_{lg},\mathcal{E}_{lg}\}$, whose feature matrix and adjacency matrix are denoted by $\mathbf{X}_{lg}$ and $\mathbf{A}_{lg}$, respectively. Specifically, the node attribute $\mathbf{X}_{lg}$ of the line graph $\mathcal{G}$ can also be defined as the edge attribute of the graph $\mathcal{G}$, i.e., $\mathbf{X}_{lg}=\mathbf{X}_{e}$. The learned node embeddings of $\mathcal{G}$ and $\mathcal{G}_{lg}$ are denoted by $\mathbf{Z}$ and $\mathbf{Z}_{lg}$, respectively. In particular, the $p$-th node in the line graph $\mathcal{G}_{lg}$ corresponds to the $p$-th connective edge between node $i$ and node $j$ in the original graph $\mathcal{G}$.
@@ IV-B Robust Edge Dropping
[…]
@@
IV-B2 GNN with Robust Edge Dropping.
@@
where $f_{\bm{\theta}}(\cdot)$ is also arbitrary GNN with parameter $\bm{\theta}$.
@@
where $\alpha$ is a trade-off hyperparameter and the node attributes of the line graph at the $k$-th epoch are updated from the learned GNN embeddings of the original graph. Namely,
$\begin{split}[\hat{\mathbf{X}}_{lg}]_{p}^{(k)}&=\frac{1}{2}\left(\operatorname{CONCAT}\left([\mathbf{Z}]_{i}^{(k)},[\mathbf{Z}]_{j}^{(k)}\right)\right.\\ &\left.+\;\operatorname{CONCAT}\left([\mathbf{Z}]_{j}^{(k)},[\mathbf{Z}]_{i}^{(k)}\right)\right).\end{split}$ (11)
Specifically, at the beginning of the training, we can initialize the node features in the line graph with a concatenation of $r(\mathbf{X}_{i})$ and $r(\mathbf{X}_{j})$, where $r(\cdot)$ is any dimensionality reduction methods such as PCA or a pretrained autoencoder. Subsequently, the model continuously updates $\mathbf{X}_{lg}$ with Eqs. (10) and (11) during training. To summarize, the edge predictor decides which edge should be removed, and the basic downstream GNN improves node attributes of the line graph $\mathcal{G}_{lg}$, i.e., input features of the edge predictor.
IV-B3 Adversarial Optimization Targets.
Then, we cope with Challenge 2 and elaborate on the targets of the network training. For the proposed framework, we need to optimize the training losses of the adversarial edge predictor and the downstream semi-supervised learning task.
@@
where $\mu$ (the same value as that in Eq. (7)) is the threshold indicating whether an edge between nodes $i$ and $j$ should be retained during training, and $\mathrm{sim}_{ij}$ is the similarity between two attributed nodes, measured by Gaussian kernel, i.e.,
@@ IV-D Complexity Analysis
Since our method is an extension of backbone GNNs, the time complexity and the memory complexity primarily depend on the architectures of the backbones. Denoting the channel number of input and output node features by $m$ and $c$, respectively, the time complexities of the compared backbones in this paper are: $\mathcal{O}(|\mathcal{E}|mc)$ (GCN), $\mathcal{O}(|\mathcal{V}|mc+|\mathcal{E}|c)$ (GAT) and $\mathcal{O}(\prod_{i=1}^{K}\mathcal{S}_{i})$ (GraphSAGE). Herein, $\mathcal{S}_{i}$ denotes the neighborhood sample size in GraphSAGE, $|\mathcal{V}|$ is the number of nodes and $|\mathcal{E}|$ is the number of edges.
@@
V Evaluation
@@ V-A1 Datasets
We consider six well-known graph datasets for performance evaluation: Citeseer [^1] https://github.com/kimiyoung/planetoid, UAI3, Pubmed1, Actor [^2] https://github.com/CUAI/Non-Homophily-Large-Scale, ACM [^3] https://github.com/zhumeiqiBUPT/AM-GCN) and Chameleon [^4] https://github.com/benedekrozemberczki/MUSAE/. The details of the graph datasets used in this paper are listed in Table I, which includes the number of nodes, edges, features, and classes. These datasets can be divided into the following types of graphs: literature citation networks (Citeseer, Pubmed and ACM, website link networks (UAI and Chameleon) and social networks (Actor). For the semi-supervised node classification task, we follow [3] and randomly adopt 20 nodes per class for training, 500 nodes for validation, and 1,000 nodes for testing.
@@
Datasets # Nodes # Edges # Features # Classes
Citeseer 3,327 4,552 3,703 6
Pubmed 19,717 44,324 500 3
ACM 3,025 13,128 1,870 3
Chameleon 2,277 31,371 2,325 5
UAI 3,067 28,308 4,973 19
Actor 7,600 26,659 932 5

### v2

@@ Abstract
Although Graph Neural Networks (GNNs) have exhibited the powerful ability to gather graph-structured information from neighborhood nodes via various message-passing mechanisms, the performance of GNNs is limited by poor generalization and fragile robustness caused by noisy and redundant graph data. As a prominent solution, Graph Augmentation Learning (GAL) has recently received increasing attention in the literature. Among the existing GAL approaches, edge-dropping methods that randomly remove edges from a graph during training are effective techniques to improve the robustness of GNNs. However, randomly dropping edges often results in bypassing critical edges. Consequently, the effectiveness of message passing is weakened. In this paper, we propose a novel adversarial edge-dropping method (ADEdgeDrop) that leverages an adversarial edge predictor guiding the removal of edges, which can be flexibly incorporated into diverse GNN backbones. Employing an adversarial training framework, the edge predictor utilizes the line graph transformed from the original graph to estimate the edges to be dropped, which improves the interpretability of the edge-dropping method. The proposed ADEdgeDrop is optimized alternately by stochastic gradient descent and projected gradient descent. Comprehensive experiments on eight graph benchmark datasets demonstrate that the proposed ADEdgeDrop outperforms state-of-the-art baselines across various GNN backbones, demonstrating improved generalization and robustness.
@@
Graph neural network, edge dropping, adversarial training, graph augmentation learning, graph representation learning.
@@ I Introduction
As powerful data representations, graphs can represent data items and their complex relations by nodes and edges, respectively. To depict node attributes and rich relational features of edges, diverse Graph Neural Networks (GNNs) have been proposed to extract distinctive node and edge representations [1, 2, 3, 4]. Representative GNN architectures include Graph Convolutional Network (GCN) [5], Graph SAmple and aggreGatE (GraphSAGE) [6], Graph Attention Network (GAT) [7], Simplifying Graph Convolution (SGC) [8], etc. During network training, GNNs propagate node features as messages, which are then passed through edges and aggregated [9]. However, inefficient message passing caused by low-quality graph data with noise and redundancy often lets GNNs suffer from overfitting, poor generalization, and less robustness [10, 11, 12].
@@
To enhance the generalization and robustness of GNN models, Graph Augmentation Learning (GAL) has emerged as a powerful technique for improving node representations through diverse data augmentation methods. Different from data augmentation in computer vision, which generates rotated or cropped images to enrich datasets, GAL focuses on learning new graph structures or updating node attributes with a diversity of strategies [13, 14]. One well-known GAL method is utilizing node feature perturbations with random corruption or an adversarial training process, thereby leading to more generalized and robust graph representations [15, 16]. Another popular approach in GALs is edge dropping, which removes or inhibits redundant edges in a graph during training to promote the robustness of GNNs. This process is analogous to the dropout technique used for training deep neural networks to mitigate overfitting. As the most commonly used edge-dropping method, the random edge-dropping strategy [17] removes edges with a certain probability, as illustrated in the left part of Figure 1. However, randomly dropping edges often results in the neglect of critical connections during training as indicated in [18, 19, 20]. Meanwhile, a graph may include noisy edges or redundant edges, which can be replaced by the integration of other multi-hop paths. Therefore, it is crucial to remove those insignificant edges while retaining those critical edges, thereby motivating us to design a new edge-dropping approach for GAL.
@@
To cope with the aforementioned issues, we propose a novel edge-dropping method, namely ADversarial Edge Drop (ADEdgeDrop) for arbitrary GNNs. Specifically, we design an adversarial edge predictor, which is responsible for determining whether to remove an edge according to the estimated edge embedding trained with a line graph derived from the original graph, as illustrated in the right part of Figure 1. To the best of our knowledge, this is the first endeavor to design a novel supervised edge-dropping method, which leverages an adversarial edge predictor to effectively remove insignificant edges while preserving critical connections. Compared with random edge-dropping methods, the proposed method has the following two advantages. (1) Better semantic interpretability owing to the edge predictor that evaluates the dropping probability of each edge through the line graph (derived from the original graph), thereby effectively capturing the relationships between the original edges. This process is supervised by the node attribute homogeneity information that evaluates the probability of correlation between two nodes. (2) Improved robustness and generalization owing to the adversarial training, consequently alleviating the overfitting effect of the edge predictor and the downstream classifier. Different from existing methods that perform adversarial training on node features, the proposed ADEdgeDrop learns a more robust graph structure by the adversarial training on edge embedding in a bottom-up manner. In summary, the technical contributions of this paper are summarized as follows.
@@
  • Comprehensive experiments validate that the proposed method obtains superior performance than existing methods with representative backbone GNNs (GCN, GAT and GraphSAGE) on eight different benchmark datasets by improving the robustness and generalization of the learned graph structure.
@@ II-A Graph Neural Networks
As a powerful tool for handling graph-structured data, GNNs have garnered significant attention, which play a crucial role in graph representation learning by effectively gathering neighborhood features and propagating vital information across the graph [21, 22, 23]. As the well-known example of GNNs, GCNs simplify the graph convolution operator on the non-Euclidean spaces by utilizing a truncated approximation of Chebyshev polynomials [5]. Additionally, Velickovic et al. defined a graph attention mechanism to explore the implicit assignment of neighborhood weights for each node [7]. Moreover, Hamilton et al. proposed GraphSAGE, a method that learns node embeddings by sampling and accumulating representations from local neighborhood nodes [6]. Furthermore, recent research attempts have been made to design simplified GNN variants to enhance GNN computations [8, 24, 25]. With the rapid development of GNNs, considerable studies have emphasized the significance of graph augmentation to enhance the GNN training and promote the robustness.
@@ II-B Graph Augmentation Learning
GAL aims to enhance the graph structure through data augmentation strategies, which are beneficial in addressing various robustness issues of GNNs [26, 27, 28]. On one hand, some endeavors have adopted adversarial training for graph representation to achieve robust node embeddings [29, 30, 31]. On the other hand, graph structure learning methods have attempted to learn richer or less connections from the original sparse graph to augment the training data. For example, Xu et al. presented a graph rewiring and preprocessing model guided by effective resistance, which can learn reasonable edge addition or removal of a graph [32]. Zhao et al. proposed the GAUG method to explore likely missing edges and remove existing edges by an edge predictor [33] though GAUG may encounter huge computational costs, particularly when dealing with a large number of nodes. As a result, edge-dropping or node-dropping strategies on existing connections are economical approaches to mitigate computational costs since recent studies have demonstrated the promising performance of these dropping methods [34], such as DropEdge [17, 35] and DropNode [36]. Additionally, Fang et al. devised a general message-dropping method that simultaneously considers various types of dropping strategies [37]. However, these methods suffer from low-quality graphs caused by random removal of nodes or edges, thereby directly influencing the performance of downstream tasks. Randomly removing edges without any criteria may have a negative impact on the graph representation, particularly if important connections are frequently eliminated during training. Hence, the key challenge of edge-dropping methods lies in identifying and removing incongruous connections rather than the essential node relationships.
@@ III-A Notations
In this paper, we define a graph as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ is the set of vertices and $\mathcal{E}$ is the set of edges. The feature matrix and the adjacency matrix of the graph $\mathcal{G}$ are denoted by $\mathbf{X}$ and $\mathbf{A}$, respectively. Similarly, the line graph transformed from the original graph $\mathcal{G}$ is represented as $\mathcal{G}_{lg}=\{\mathcal{V}_{lg},\mathcal{E}_{lg}\}$, whose feature matrix and adjacency matrix are denoted by $\mathbf{X}_{lg}$ and $\mathbf{A}_{lg}$, respectively. Specifically, the node attribute $\mathbf{X}_{lg}$ of the line graph $\mathcal{G}_{lg}$ can also be defined as the edge attribute of the graph $\mathcal{G}$, i.e., $\mathbf{X}_{lg}=\mathbf{X}_{e}$. The learned node embeddings of $\mathcal{G}$ and $\mathcal{G}_{lg}$ are denoted by $\mathbf{Z}$ and $\mathbf{Z}_{lg}$, respectively. In particular, the $p$-th node in the line graph $\mathcal{G}_{lg}$ corresponds to the $p$-th connective edge between node $i$ and node $j$ in the original graph $\mathcal{G}$.
@@ IV-B1 Edge Predictor
$\left[\mathbf{A}_{lg}\right]_{(i\to j),(i^{\prime}\to j^{\prime})}=\begin{cases}1&j=i^{\prime},i\neq j^{\prime},\\ 0&\mathrm{otherwise}.\end{cases}$ (4)
@@
IV-B2 GNN with Robust Edge Dropping
@@
where $f_{\bm{\theta}}(\cdot)$ is also an arbitrary GNN with parameter $\bm{\theta}$.
@@
where $\alpha$ is a trade-off hyperparameter, and the node attributes of the line graph at the $k$-th epoch are updated from the learned GNN embeddings of the original graph, namely,
$[\hat{\mathbf{X}}_{lg}]_{p}^{(k)}=\mathrm{softmax}\left(\operatorname{CONCAT}\left([\mathbf{Z}]_{i}^{(k)},[\mathbf{Z}]_{j}^{(k)}\right)\right..$ (11)
Specifically, at the beginning of the training, we can initialize the node features in the line graph with a concatenation of $r(\mathbf{X}_{i})$ and $r(\mathbf{X}_{j})$, where $r(\cdot)$ is any dimensionality reduction method such as PCA or a pretrained autoencoder. Subsequently, the model continuously updates $\mathbf{X}_{lg}$ with Eqs. (10) and (11) during training. To summarize, the edge predictor decides which edge should be removed, and the basic downstream GNN improves node attributes of the line graph $\mathcal{G}_{lg}$, i.e., input features of the edge predictor.
IV-B3 Adversarial Optimization Targets
We cope with Challenge 2 and elaborate on the targets of the network training. For the proposed framework, we need to optimize the training losses of the adversarial edge predictor and the downstream semi-supervised learning task.
@@ IV-B3 Adversarial Optimization Targets
where $\mu$ (the same value as that in Eq. (7)) is the threshold indicating whether an edge between nodes $i$ and $j$ should be retained during training, and $\mathrm{sim}_{ij}$ is the similarity between two attributed nodes, measured by the Gaussian kernel, i.e.,
@@ IV-D Complexity Analysis
Since our method is an extension of backbone GNNs, the time complexity and the memory complexity primarily depend on the architectures of the backbones. Herein, we take the applied backbones in our experiments as examples, i.e., GCN [5], GAT [7] and GraphSAGE [6]. Denoting the channel number of input and output node features by $m$ and $c$, respectively, the time complexities of the compared backbones in this paper are: $\mathcal{O}(|\mathcal{E}|mc)$ (GCN), $\mathcal{O}(|\mathcal{V}|mc+|\mathcal{E}|c)$ (GAT) and $\mathcal{O}(\prod_{i=1}^{K}\mathcal{S}_{i})$ (GraphSAGE). Herein, $\mathcal{S}_{i}$ denotes the neighborhood sample size in GraphSAGE, $|\mathcal{V}|$ is the number of nodes and $|\mathcal{E}|$ is the number of edges.
@@
IV-E Strategy for Edge-Dense Graphs
From the previous analyses, it can be found that the time consumption of most edge-dropping methods is tightly related to the number of edges in the graph. In order to cope with large-scale datasets, especially graphs with a large number of edges, we further propose a strategy for the edge-dropping strategy with graphs where edges are dense. The proposed method conducts a pre-dropping process on the line graph to reduce the computational cost of the edge predictor on line graphs. Because we assume that the dropping strategy of important edges that connect semantically similar nodes should be seriously considered, our strategy for edge-dense graphs only constructs line graphs on these critical edges. Namely, we conduct the pre-dropping process according to node semantic similarities by
$\hat{\mathbf{A}}_{ij}=\mathbf{M}_{ij}\odot\mathbf{A}_{ij},$ (20)
where $\mathbf{M}_{ij}$ is the estimated edge mask computed by
$\mathbf{M}_{ij}=\begin{cases}1&\mathrm{sim}_{ij}\geq p_{pre},\\ 0&\mathrm{sim}_{ij}<p_{pre},\end{cases}$ (21)
where $p_{pre}$ is the pre-dropping rate, and $\mathrm{sim}_{ij}$ has been pre-computed by Eq. (14). During GNN training, we also conduct an ordinary edge dropping on the original graph to remove more edges randomly. Nevertheless, in contrast to the sole edge-dropping methods, the proposed method preserves critical edges during training with the learnable adversarial edge removal w.r.t. semantic-important node relationships.
IV-F Discussion
The proposed ADEdgeDrop aims to perform a more robust edge-dropping procedure during GNN training. They differ from the existing methods in the following aspects:
  • ADEdgeDrop is a trainable edge-dropping procedure, which adaptively adjusts the removed edges during different GNN training procedures according to node semantic similarities with an adversarial optimization objective.
  • Different from existing adversarial-training-based methods (e.g., FLAG [16]) that learn robust node embeddings, the proposed method conducts the adversarial training on the line graphs including the relationships between edges in the original graphs, thereby enabling the adversarial training w.r.t. edge embeddings in the original graphs.
  • Consequently, the proposed method is an adversarial edge-dropping strategy, which can improve the robustness of distinct GNNs and offer insights for future GNN training on a sparser graph with fewer edges.
V Experimental Evaluation
@@ 3. RQ3:
It is noted that our experiments not only focus on performance and training benefits obtained from the proposed edge-dropping method, but also concentrate on the learned improved graph structure with sparser edges.
@@ V-A1 Datasets
We consider eight well-known graph datasets for performance evaluation: BlogCatalog [^1] https://networkrepository.com/soc-BlogCatalog.php, Pubmed [^2] https://github.com/kimiyoung/planetoid, ACM [^3] https://github.com/zhumeiqiBUPT/AM-GCN, Chameleon [^4] https://github.com/benedekrozemberczki/MUSAE/, UAI3, Actor [^5] https://github.com/CUAI/Non-Homophily-Large-Scale, CoraFull [^6] https://github.com/shchur/gnn-benchmark#datasets and Arxiv [^7] https://ogb.stanford.edu/docs/nodeprop/. The details of the graph datasets used in this paper are listed in Table I, which summarizes the number of nodes, edges, features, and classes. For the semi-supervised node classification task, we follow [5] and randomly adopt 20 nodes per class for training, 500 nodes for validation, and 1,000 nodes for testing. Specifically, we follow the publicly available split mode for the large-scale Arxiv dataset.
@@
Datasets BlogCatalog Pubmed ACM Chameleon UAI Actor CoraFull Arxiv
# Nodes 5,196 19,717 3,025 2,277 3,067 7,600 19,793 169,343
# Edges 141,738 44,324 13,128 31,371 28,308 26,659 63,421 1,157,799
# Features 8,189 500 1,870 2,325 4,973 932 8,710 128
# Classes 6 3 3 5 19 5 70 40